# Debating Truth: Debate-driven Claim Verification with Multiple Large Language Model Agents

### v2

@@ 5.5. Impact of LLMs for Debaters
5.6. Computational Cost Analysis
Table 6. Computational cost per claim verification.
Role Input Output
Total
Cost (\$)
Tokens Cost (\$) Tokens Cost (\$)
Debaters (GPT-4o-Mini) 3685.13 0.0005 1526.01 0.0009 0.0014
Moderator (Qwen-2.5-7B) 3355.58 0.0006 290.78 0.0002 0.0008
Total 7040.71 0.0011 1816.79 0.0011 0.0022
As shown in Table 6, verifying one claim costs \$0.0014 for Debaters and \$0.0008 for the Moderator, totaling \$0.0022. For context, PolitiFact published 74 fact-check reports from July 1 to September 30, 2024, as counted on their official website. Using DebateCV would cost $74\times 0.0022=\mathdollar 0.1628$, which can be significantly more cost-efficient than human fact-checkers. While we acknowledge that DebateCV can be relatively more expensive than single-agent methods like HerO, this additional cost is common in multi-agent frameworks (debate_improve_factuality; mad; lan2024stance) and is justified by accuracy gains, especially in high-stakes applications like fact-checking, where even marginal improvements can yield substantial real-world impact.
@@ 6. Conclusion
In this work, we introduce DebateCV, the first debate-driven claim verification framework. DebateCV orchestrates multiple LLM agents in structured adversarial debates, where two Debaters take opposing stances and present evidence-grounded arguments, while a Moderator synthesises these debates into final verdicts. To improve the Moderator’s ability to adjudicate complex, multi-round debates, we propose Debate-SFT, a novel post-training framework that leverages synthetic debate data generated through a three-stage pipeline: debate data synthesis, error correction, and post-training. Extensive experiments on the state-of-the-art AVeriTeC benchmark show that DebateCV significantly outperforms both single-agent and multi-agent baselines across a range of evidence quality conditions. Our analysis reveals that these performance gains stem from DebateCV’s adversarial cross-validation mechanism to mitigate three fundamental error patterns prevalent in existing methods: claim/evidence misinterpretation, overlooking relevant evidence, and overreliance on speculation. Beyond quantitative improvements, human expert evaluations show that DebateCV produces substantially higher-quality justifications, enhancing the transparency and trustworthiness essential for practical fact-checking applications. These technical advancements will enhance digital literacy and strengthen misinformation countermeasures with transparent justifications, thereby fostering a more trustworthy information ecosystem in alignment with UN Sustainable Development Goals, such as good health and well-being (Goal 3) and sustainable cities and communities (Goal 11).
@@ Appendix A Definitions of the Verdict Categories
Similar to (averitec; infact; hero; CTU_at_averitec), we consider four categories of verdicts for the claims. Below is a summary of their definitions. (i) Supported: The claim is supported by the arguments and evidence presented. (ii) Refuted: The claim is contradicted by the arguments and evidence presented. (iii) Not Enough Evidence: The presented evidence is not enough to support or refute the claim. This category applies when the evidence either explicitly indicates that relevant evidence cannot be found or leaves certain aspects of the claim neither supported nor refuted. (iv) Conflicting Evidence/Cherry-picking: The claim is misleading due to conflicting evidence or cherry-picking, but is not explicitly refuted. This category includes cases such as cherry-picking (selectively presenting evidence to misrepresent truth), true-but-misleading (e.g., “Alice has never lost an election” when Alice has only ever run unopposed), and instances where contradictory evidence is found. A more detailed description of these categories and additional examples can be found in (averitec).
@@ B.1. Pre-debate Configuration Prompts
- Address and critique the opposing evidence to strengthen your position, while being concise.
@@ B.2. In-debate Process Prompts
Table 7. Examples of DebateCV outputs for claim verification. The accurate judgments are in green.
@@
UNESCO declared Nadar community as the most ancient race in the world. (ID: 3; Refuted) The Nadar community has a rich historical presence and cultural significance. Some scholars recognize their deep roots. Supported. No direct recognition from UNESCO supporting this claim. Many communities have ancient lineages, and cultural prominence does not equate to being the “most ancient.” Refuted. While the Affirmative highlights historical importance, it lacks official UNESCO endorsement. The Negative argument effectively refutes the lack of formal recognition. Refuted.
@@
The United States and its Western allies have been using their media outlets to publish articles based on fabricated information under allegations of non-compliance with the Chemical Weapons Convention. (ID: 5; Refuted) There are accusations from multiple sources—including pro-Kremlin and Western outlets—that narratives around Syria’s compliance were manipulated. OPCW reports have also been contested. Supported. OPCW investigations consistently raised concerns about Syria’s compliance, supported by verified findings. Reports of chemical use are corroborated by independent missions. Allegations are based on real evidence. Refuted. The Negative side provided stronger and more concrete evidence, including verified OPCW findings. Affirmative claims lacked proof of systemic fabrication. Refuted.
@@
“Why should you pay more taxes than Donald Trump pays?” Biden cited \$750 paid by Trump in 2016 and 2017, and quoted Trump saying, “Because I’m smart. I know how to game the system.” (ID: 7; Supported) Multiple reputable sources (NYT, NPR, Forbes) confirm Trump paid only \$750 in federal income tax during those years. His own quote suggests he exploited legal loopholes. Supported. Breitbart disputes the NYT report, suggesting it misrepresents Trump’s filings. Not all wealthy individuals avoid taxes, but many legally reduce liabilities through established strategies. The claim oversimplifies tax complexities. Refuted. Despite counterarguments, the Affirmative side presented credible and consistent evidence from multiple reputable sources. The Negative side failed to refute the core facts. Supported.
@@
Joe Biden wants to ban fracking. (ID: 23; Refuted) Biden supports banning new oil and gas permits on public lands, which limits fracking. While he hasn’t explicitly called for a ban, his campaign indicated restrictions that align with a de facto ban. Supported. Biden has explicitly stated he does not support a complete ban on fracking. He aims to restrict new permits on public lands, not eliminate existing operations. Campaign statements clarify this nuance. Refuted. The Negative correctly identifies Biden’s nuanced stance. His policies target new permits, not an outright ban. Refuted.
@@
The New York Post endorsed Donald Trump’s bid for re-election. (ID: 32; Supported) The New York Post editorial board explicitly endorsed Trump in the 2020 election, despite some policy criticisms. Endorsements can reflect strategic alignment rather than full agreement. Supported. While the Post has conservative leanings, its endorsement included reservations and critiques, suggesting a complex rather than clear-cut stance. Refuted. The Affirmative presents direct evidence of the endorsement, while the Negative raises contextual concerns without disproving it. Supported.
@@
The COVID-19 pandemic was pre-planned with the help of Bill Gates. (ID: 35; Refuted) Gates has long warned of pandemics. These warnings suggest foreknowledge and planning, fueling public perception of involvement. Supported. Gates’ advocacy was for preparedness, not prediction. Event 201 was a simulation, not a blueprint. Conspiracy theories lack factual basis. Refuted. Affirmative relies on correlation and speculation, while the Negative clearly distinguishes between warning and orchestration. Refuted.
@@
Breitbart News reports that the daughter of Delaware Democratic Senator Chris Coons and seven other underage girls were featured on Hunter Biden’s laptop. (ID: 37; Refuted) Allegations from political figures like Lauren Witzke suggest there may be merit to the claim. Lack of denial implies some level of credibility. Supported. No credible source confirms Breitbart reported this. Fact-checks disprove child pornography claims. Unsupported allegations should not be treated as fact. Refuted. Affirmative relies on unverified allegations, while the Negative shows absence of actual reporting or evidence. Refuted.
@@
The umbilical cord should not be cut until 1 hour after birth or the baby will not have enough blood right after birth. (ID: 46; Refuted) ACOG recommends delayed clamping (30–60 sec) to improve hemoglobin and iron levels. A 1-hour delay may offer additional benefits. Supported. ACOG supports only 30–60 seconds of delay. Longer delay lacks evidence and may not be practical. Refuted. Delayed clamping is beneficial, but notes no support for 1-hour delay. Medical guidelines back shorter window. Refuted.
@@
People who do not vote for the Bharatiya Janata Party (BJP) in the 2020 elections will not get the COVID vaccine free of cost. (ID: 51; Refuted) BJP manifesto promised free vaccines, but critics feared this might be tied to voter loyalty. Political context implied exclusion. Supported. Official statements clarified vaccines would be available to all priority groups regardless of political affiliation. Refuted. No evidence linking eligibility to voting behavior. Government immunization programs apply universally. Refuted.
@@
Officer who wore Trump 2020 mask to polls to face disciplinary action. (ID: 63; Supported) A Miami police officer was photographed wearing a ’Trump 2020’ mask at a polling location. NBC News and Miami PD confirm disciplinary action is coming. Mayor Suarez affirmed consequences. Supported. While the officer likely wore the mask and may face discipline, the term ’disciplinary action’ is vague. Free speech rights and political expression complicate the issue. Refuted. Multiple evidence confirms both the mask-wearing and disciplinary action. The Negative raises valid questions but doesn’t contradict the central claim. Supported.
@@
1. Summarize the main new insights obtained from this round compared to previous rounds.
@@ 4. Conclusion:
Opposition Argument: [OPPOSITION ARGUMENT] Do you agree with this perspective? Provide your response explaining your reasoning, supporting evidence, and highlighting any weaknesses in the opposition’s points.
If the Moderator determines that the debate between the two Debaters has converged, the debate will be terminated. In cases where the debate lasts for more than three rounds, the Moderator receives the following final prompt to output a verdict and generate a corresponding justification.
@@ B.3. Corrector’s Prompt
This prompt instructs the Corrector to generate a justification explaining how the ground-truth verdict can be derived from the debate.
Debate Recording: [DEBATE_RECORDING] Primary Insights: [Primary_Insights] Task: Please provide the justification for your verdict that this claim is [GT_VERDICT] based on the debate context. The output should be in JSON format. { "Justification for Verdict": "…" }
@@ Appendix C Examples of DebateCV Outputs
To demonstrate how DebateCV performs claim verification, we provide ten examples in Table 7, in addition to the case study detailed in Section 5. Each example consists of an input claim, followed by arguments presented by the Affirmative and Negative Debaters, and concludes with the Moderator’s final decision. We use the following prompt with GPT-4o-Mini to condense agent responses across all debate rounds for brevity:
This is a claim verification debate that involves two Debaters (Affirmative and Negative) and a Moderator. The claim is: [CLAIM] The full transcript of the debate is as follows: [DEBATE_RECORD] Summarize the arguments presented by both sides and the Moderator’s final decision. Write each summary in the first person from the perspective of each Debater and the Moderator.

### v3

@@ 6. Conclusion
In this work, we introduce DebateCV, the first debate-driven claim verification framework. DebateCV orchestrates multiple LLM agents in structured adversarial debates, where two Debaters take opposing stances and present evidence-grounded arguments, while a Moderator synthesises these debates into final verdicts. To improve the Moderator’s ability to adjudicate complex, multi-round debates, we propose Debate-SFT, a novel post-training framework that leverages synthetic debate data generated through a three-stage pipeline: debate data synthesis, error correction, and post-training. Extensive experiments on the state-of-the-art AVeriTeC benchmark show that DebateCV significantly outperforms both single-agent and multi-agent baselines across a range of evidence quality conditions. Our analysis reveals that these performance gains stem from DebateCV’s adversarial cross-validation mechanism to mitigate three fundamental error patterns prevalent in existing methods: claim/evidence misinterpretation, overlooking relevant evidence, and overreliance on speculation. Beyond quantitative improvements, human expert evaluations show that DebateCV produces substantially higher-quality justifications, enhancing the transparency and trustworthiness essential for practical fact-checking applications. These technical advancements will enhance digital literacy and strengthen misinformation countermeasures with high-quality justifications, thereby fostering a more trustworthy information ecosystem. Future research could explore methods to enhance the inference cost efficiency of DebateCV, [^16] Appendix D analysed the inference cost of DebateCV. as well as strategies to improve claim verification performance under extremely noisy or even heavily polluted evidence conditions.
Acknowledgments
This work was done while Haorui He was under the supervision of Yupeng Li. This work was supported by National Natural Science Foundation of China (No. 62202402), Guangdong and Hong Kong Universities “1+1+1” Joint Research Collaboration Scheme, Project No. 2025A0505000001, the Early Career Scheme (ECS) from the Research Grants Council of HKSAR (HKBU 22202423), the General Research Fund (GRF) from the Research Grants Council of HKSAR (HKBU 12203425), a grant from the Germany/Hong Kong Joint Research Scheme sponsored by the Research Grants Council of HKSAR and the German Academic Exchange Service of Germany (No. G-HKBU208/25), the Initiation Grant for Faculty Niche Research Areas 2023/24 (No. RC-FNRA-IG/23-24/COMM/01), Research Cluster Matching Scheme (No. RCMS/24-25/01) of Hong Kong Baptist University, National Natural Science Foundation of China (No. 62072115), the Shanghai Science and Technology Innovation Action Plan Project (No. 22510713600), the grants from the Research Grants Council of HKSAR (HKU 17202325), the University of Hong Kong (Project 2409100399), the HKU Faculty Exchange Award 2024 (Faculty of Engineering), Scientific Research Innovation Capability Support Project for Young Faculty (No. SRICSPYF-BS2025137), and Startup Grant (Tier 1) for New Academics AY2020/21 of Hong Kong Baptist University.
@@ Appendix A Definitions of the Verdict Categories
Similar to (Schlichtkrull et al., 2024b; Rothermel et al., 2024; Yoon et al., 2024), we consider four categories of verdicts for the claims. Below is a summary of their definitions. (i) Supported: The claim is supported by the arguments and evidence presented. (ii) Refuted: The claim is contradicted by the arguments and evidence presented. (iii) Not Enough Evidence: The presented evidence is not enough to support or refute the claim. This category applies when the evidence either explicitly indicates that relevant evidence cannot be found or leaves certain aspects of the claim neither supported nor refuted. (iv) Conflicting Evidence/Cherry-picking: The claim is misleading due to conflicting evidence or cherry-picking, but is not explicitly refuted. This category includes cases such as cherry-picking (selectively presenting evidence to misrepresent truth), true-but-misleading (e.g., “Alice has never lost an election” when Alice has only ever run unopposed), and instances where contradictory evidence is found. Please see (Schlichtkrull et al., 2024b) for a detailed description of these categories and examples.
@@ B.1. Pre-debate Configuration Prompts
- Critique the opposing evidence to strengthen your position, while being concise.
@@ B.2. In-debate Process Prompts
Table 6. Examples of DebateCV outputs for claim verification. Accurate judgments are in green.
@@
UNESCO declared Nadar community as the most ancient race in the world. (ID: 3; Refuted) The Nadar community has a rich historical presence and cultural significance. Some scholars recognize their deep roots. Supported. No direct recognition is from UNESCO. Many communities have ancient lineages, and cultural prominence does not equate to “most ancient.” Refuted. The Negative effectively refutes the Affirmative by highlighting its lack of official UNESCO endorsement. Refuted.
@@
The USA and its Western allies have been using their media outlets to publish articles based on fabricated information under allegations of non-compliance with the Chemical Weapons Convention. (ID: 5; Refuted) There are accusations from multiple sources—including pro-Kremlin and Western outlets—that narratives around Syria’s compliance were manipulated. OPCW reports have also been contested. Supported. OPCW investigations consistently raised concerns about Syria’s compliance, supported by verified findings. Reports of chemical use are corroborated by independent missions. Allegations are based on real evidence. Refuted. The Negative side provided stronger and more concrete evidence, including verified OPCW findings. Affirmative claims lacked proof of systemic fabrication. Refuted.
@@
“Why should you pay more taxes than Donald Trump pays?” Biden cited \$750 paid by Trump in 2016 and 2017, and quoted Trump saying, “Because I’m smart. I know how to game the system.” (ID: 7; Supported) Multiple reputable sources (NYT, NPR, Forbes) confirm Trump paid only \$750 in federal income tax during those years. His own quote suggests he exploited legal loopholes. Supported. Breitbart disputes the report, suggesting it misrepresents Trump’s filings. Not all wealthy individuals avoid taxes, but many legally reduce liabilities through established strategies. The claim oversimplifies tax complexities. Refuted. Despite counterarguments, the Affirmative side presented credible and consistent evidence from multiple reputable sources. The Negative side failed to refute the core facts. Supported.
@@
Joe Biden wants to ban fracking. (ID: 23; Refuted) Biden supports banning new oil and gas permits on public lands, which limits fracking. While he hasn’t explicitly called for a ban, his campaign indicated restrictions that align with a de facto ban. Supported. Biden explicitly stated he does not support a complete ban, instead he aims to restrict new permits on public lands. Campaign statements clarify this nuance. Refuted. The Negative correctly identifies Biden’s nuanced stance. His policies target new permits, not an outright ban. Refuted.
@@
The New York Post endorsed Donald Trump’s bid for re-election. (ID: 32; Supported) The NYP editorial board explicitly endorsed Trump in the 2020 election, despite some policy criticisms, which reflects strategic alignment rather than full agreement. Supported. While the Post has conservative leanings, its endorsement included reservations and critiques, suggesting a complex rather than clear-cut stance. Refuted. The Affirmative presents direct evidence of the endorsement, while the Negative raises contextual concerns without disproving it. Supported.
@@
The COVID-19 pandemic was pre-planned with the help of Bill Gates. (ID: 35; Refuted) Gates has long warned of pandemics, which suggests foreknowledge and planning, fueling public perception of involvement. Supported. The advocacy was for preparedness, not prediction. Conspiracy theories lack factual basis. Refuted. Affirmative relies on speculation, while the Negative clearly distinguishes between warning and orchestration. Refuted.
@@
Breitbart News reports that the daughter of Delaware Democratic Senator Chris and seven other underage girls were featured on Hunter Biden’s laptop. (ID: 37; Refuted) Allegations from political figures like Lauren Witzke suggest there may be merit to the claim. Lack of denial implies some level of credibility. Supported. No credible source confirms Breitbart reported this. Fact-checks disprove child pornography claims. Unsupported allegations should not be treated as fact. Refuted. Affirmative relies on unverified allegations, while the Negative shows absence of actual reporting or evidence. Refuted.
@@
The umbilical cord should not be cut until 1 hour after birth or the baby will not have enough blood right (ID: 46; Refuted) ACOG recommends delayed clamping (30–60 sec) to improve hemoglobin and iron levels. A 1-hour delay may offer benefits. Supported. ACOG supports only 30–60 seconds of delay. Longer delay lacks evidence and may not be practical. Refuted. Delayed clamping is beneficial, but notes no support for 1-hour delay. Medical guidelines back shorter window. Refuted.
@@
People who do not vote for the BJP in the 2020 elections will not get the COVID vaccine free of charge. (ID: 51; Refuted) BJP manifesto promised free vaccines, but critics feared this might be tied to voter loyalty. Political context implied exclusion. Supported. Official statements clarified vaccines would be available to all priority groups regardless of political affiliation. Refuted. No evidence linking eligibility to voting behavior. Government immunization programs apply universally. Refuted.
@@
Officer who wore Trump 2020 mask to polls to face disciplinary action. (ID: 63; Supported) A police officer was photographed wearing a ’Trump 2020’ mask at a polling location. Mayor Suarez affirmed consequences. Supported. While the officer likely wore the mask and may face discipline, the term ’disciplinary action’ is vague. Free speech rights and political expression complicate the issue. Refuted. Multiple evidences confirm both the mask-wearing and disciplinary action. The Negative raises valid questions but doesn’t contradict the central claim. Supported.
@@
1. Summarize the new insights from this round compared to previous rounds.
@@ 4. Conclusion:
Opposition Argument: [OPPOSITION ARGUMENT] Do you agree with this perspective? Provide your response explaining your reasoning, supporting evidence, and highlighting any weaknesses in the opposition.
If the Moderator determines that the debate has converged, the debate will be terminated. In cases where the debate lasts for more than three rounds, the Moderator receives the final prompt to output a verdict and generate a corresponding justification.
@@ B.3. Corrector’s Prompt
Corrector generates a justification explaining how the ground-truth verdict can be derived from the debate.
Debate Recording: [DEBATE_RECORDING] Primary Insights: [Primary_Insights] Task: Please provide the justification for the verdict that this claim is [GT_VERDICT] based on the debate context. The output should be in JSON format. { "Justification for Verdict": "…" }
@@ Appendix C Examples of DebateCV Outputs
To demonstrate how DebateCV performs claim verification, we provide ten examples in Table 6, in addition to the case study detailed in Section 5. Each example consists of an input claim, followed by arguments presented by the Affirmative and Negative Debaters, and concludes with the Moderator’s final decision. Agent responses are condensed across all debate rounds for brevity.
Appendix D Computational Cost Analysis
Table 7. Computational cost per claim verification.
Role Input Output Total Cost (\$)
Tokens Cost (\$) Tokens Cost (\$)
Debaters (GPT-4o-Mini) 3685.13 0.0005 1526.01 0.0009 0.0014
Moderator (Qwen-2.5-7B) 3355.58 0.0006 290.78 0.0002 0.0008
Total 7040.71 0.0011 1816.79 0.0011 0.0022
As shown in Table 7, verifying one claim costs \$0.0014 for Debaters and \$0.0008 for the Moderator, totaling \$0.0022. For context, PolitiFact published 74 fact-check reports from July 1 to September 30, 2024, as counted on their official website. Using DebateCV would cost $74\times 0.0022=\mathdollar 0.1628$, which is obviously more cost-efficient than human fact-checkers. While we acknowledge that DebateCV can be relatively more expensive than single-agent methods like HerO, this additional cost is common in multi-agent frameworks (Du et al., 2024; Liang et al., 2024) and is justified by the significant accuracy gains, especially in high-stakes applications like fact-checking, where even marginal improvements can yield substantial real-world impact.